Telescope: Clay
Instrument: MIKE
Observer(s): Aarnio, Irwin
Assistant: Sergio

File	   Object		UT start	Exp.[s]	Airmass	Slit	Binning	Comments
148	   ThAr			19:59		15		0".35	1x1	Fast readout mode (86s)
149-158	   Quartz		20:06		120/30		"	"
159	   ThAr			20:42		15		
160-169	   Milky Quartz		20:46		120/96				Diffuser in
170-179	   Bias	 		21:22		0				Diffuser out
180-189	   Bias			21:37		0		1"	2x2	slow readout; begin Irwin calibrators
190	   ThAr			22:32		5			
191-200	   Quartz		22:33		6/3			
201	   ThAr			22:44		5
202-211	   Milky Quartz		22:46		20/16				Diffuser in
212	   ThAr	 		23:03		5
213	   ThAr			23:29		5
214-223	   Twilight trace	23:31		10		1"	2x2
224	   ThAr	    		23:43		5
225-234	   1331852		23:59		5/14				V=2.2, rapid rotator. BAD!! Diffuser was in part-way!!!
235	   ThAr			00:11		5				Diffuser out
236	   ThAr			00:13		15		0".35	1x1	
237-246	   1331852		00:16		67/36	1.146			Diffuser in. BAD!!! Diffuser stuck partway!!
247	   ThAr			00:45		15				BAD!!! Diffuser still stuck partway!!
248	   AB Aur		00:47		280/160				BAD!! Diffuser partway in [we discovered problem on this object]
249	   AB Aur		00:56		5				testing whether issue is resolved, no it's not
250	   AB Aur		01:00		5				testing again whether issue is resolved- yes it is
251	   AB Aur		01:03		280/160	2.082			
252	   AB Aur				320/200
253	   AB Aur				350/200
254-255	   ThAr			01:24		15
256	   T Tau		01:28		550/250	1.801			Seeing ~0".7, humidity 63%
257	   T Tau		01:40		1000/700 1.878			Setting fast, low cts even at 1000/700s [we already have MIKE data for this star...]
258-259	   Th Ar		02:08		15
260-262	   HR 1544		02:28		30/25	1.57			A1 IV flux calibrator
263-264	   ThAr			02:34		15
265	   VY Mon		02:38		840/404	1.347			abysmally low counts on blue side
266	   VY Mon		03:06		2000/1500 1.399			Even 33 minutes didn't help the blue side. Not going to spend any more time on this object
267-268	   ThAr			03:42		15
269-271	   HD 50138		03:47		250/155	1.200
272-273	   ThAr			04:04		15
274-276	   HD 60753		04:09		200/150	1.131
277	   ThAr			04:27		15
278-280	   MWC 166		04:29		250/160	1.252			B3 IV flux calibrator
281-282	   ThAr			04:46		15
283-285	   HD 58647		04:50		220/150	1.214
286-287	   ThAr			05:07		15
288-290	   HR 3454		05:20		15/20	1.261
291-292	   ThAr			05:26		15
293	   TW Hya		05:32		550/315	1.016
294-295	   TW Hya		05:44		650/400	1.009			Whopping emission lines on red side!
296-297	   ThAr			06:09		15
298	   Procyon		06:14		2/9	2.054			Defocused at 100, elevation=30 deg. Both sides saturated
299	   Procyon		06:17		4/4				red side still saturated
300	   Procyon		06:19		4/2	2.080			red side still saturated
301-303	   Procyon		06:21		4/1	2.104
304-305	   ThAr			06:26		15
Begin 2.5 hours for Irwin
End of night calibrations (rapid rotator again)
332    	   ThAr		  	09:32		15
333-342	   1895685		09:34		94/52	1.083			V=2.6 rapid rotator. Diffuser in
343	   ThAr			10:06		15				Diffuser out
END OF NIGHT